An unknown charged particle passes without deflect
Answer #1

A. using force balance

Fe = Fm

qE = qVB

E = VB

V = E/B = 187500/0.125

V = 1.5*10^6 m

Now using circular motion

Fc = Fm

mV^2/r = qVB

q/m = V/rB

r = d/2 = 0.2505/2

q/m = 1.5*10^6*2/(0.2505*0.125)

q/m = 9.58*10^7 C/kg

B.

for proton

q/m = 1.6*10^-19/(1.67*10^-27) =9.58*10^7

for electron

q/m = 1.6*10^-19/(9.1*10^-31) = 1.758*10^11

for neutron

q/m = 0/(1.67*10^-24) = 0

So correct answer will be proton.